a.Why are the lines whose equations are y+5 and y3x 1 perpendicular? b.Use a graphing utility to graph the equations in a- 10,10,1] by 10,10,1] viewing rectangle. Do the lines appear to be perpendicular? c. Now use the zoom square feature of your utility. Describe what happens to the graphs. Explain why this is so.
Answer #1

a) The lines

3 y=-x + 5 and  y=-3x-1 are perpendicular to each other since the product of their slopes are -1.

That is (1/3)*(-3)=-1

b) By using Graphing Utility:

From the graph of the lines of the given equations it is clear that the two lines are perpendicular.

c) By using Zoom square feature the angle between the intersectionof two lines seems to be exactly 900 since the precision of measurment increases.